OpenStudy (anonymous):

Given the equation y − 4 =3/4 (x + 8) in point-slope form, identify the equation of the same line in standard form

OpenStudy (ash2326):

standard form y=mx+c m-slope c-y intercept y-4=3/4(x+8) y=3/4 x+6+4 y=3/4 x +10 slope 3/4 intercept 10